### v1

# AD-R1: Closed-Loop Reinforcement Learning for End-to-End Autonomous Driving with Impartial World Models

Abstract
End-to-end models for autonomous driving hold the promise of learning complex behaviors directly from sensor data, but face critical challenges in safety and handling long-tail events. Reinforcement Learning (RL) offers a promising path to overcome these limitations, yet its success in autonomous driving has been elusive. We identify a fundamental flaw hindering this progress: a deep-seated “optimistic bias” in the world models used for RL. Trained exclusively on safe expert data, these models fail to predict the consequences of erroneous actions. When conditioned on an unsafe trajectory, they hallucinate an implausibly safe future—obstacles vanish, roads are ignored—rather than forecasting the impending failure. This inability to “imagine failure” makes them unreliable critics for learning robust driving policies. To address this, we introduce a framework for post-training policy refinement built around an Impartial World Model. Our primary contribution is to teach this model to be honest about danger. We achieve this with a novel data synthesis pipeline, Counterfactual Synthesis, which systematically generates a rich curriculum of plausible collisions and off-road events. This transforms the model from a passive scene completer into a veridical forecaster that remains faithful to the causal link between actions and outcomes. We then integrate this Impartial World Model into our closed-loop RL framework, where it serves as an internal critic. During refinement, the agent queries the critic to “dream” of the outcomes for candidate actions. We demonstrate through extensive experiments, including on a new Risk Foreseeing Benchmark, that our model significantly outperforms baselines in predicting failures. Consequently, when used as a critic, it enables a substantial reduction in safety violations in challenging simulations, proving that teaching a model to dream of danger is a critical step towards building truly safe and intelligent autonomous agents.
1 Introduction
Figure 1: (a) Conventional RL Post-training relies on external simulators, suffering from a sim-to-real gap and heuristic rewards. (b) Our Paradigm uses a learned Occupancy World Model as an internal, generative simulator, enabling learned interactions and physically-grounded rewards. (c) The “Optimistic Bias”: standard world models, however, trained only on safe data, fail to predict danger. They hallucinate safe futures for unsafe plans, providing dangerously high rewards. (d) Our Impartial World Model, trained with Counterfactual Synthesis Data, learns to imagine failure. It faithfully predicts the hazardous outcome, providing a correct, punitive reward, and enabling the agent to learn safely.
End-to-end autonomous driving [hu2023uniad, jiang2023vad, chen2024vadv2, li2025recogdrive, wote, law, liao2025diffusiondrive, xing2025goalflow] has recently revolutionized the field, but policies trained with Imitation Learning (IL) still face significant challenges in real-world deployment, such as failures in long-tail events due to distributional shift [gao2025rad, li2025recogdrive]. Consequently, the field is increasingly shifting towards closed-loop RL [guo2025deepseek, ouyang2022rlhf, lu2025vla-rl] through active environmental interaction to enhance decision-making capabilities of agents.
However, the path to applying RL effectively in autonomous driving is fraught with challenges, and its widespread success has remained elusive. A fundamental obstacle stems from the very environment where RL agents are typically trained. Conventional approaches [li2024think2drive, yang2025raw2drive, li2025recogdrive] must rely on external simulators [dauner2024navsim, dosovitskiy2017carla], which introduce their own set of limitations. These include the critical sim-to-real gap inherent in graphics engines [dosovitskiy2017carla, li2024think2drive, yang2025raw2drive], the non-interactive nature and lack of 3D geometric awareness in traffic-flow simulators [dauner2024navsim, li2025recogdrive, liu2025reinforced]. These persistent issues create a major bottleneck, limiting the transferability and effectiveness of RL-trained policies in the real world. These persistent issues necessitate a paradigm shift. Inspired by the remarkable success of recent generative models [gao2023magicdrive, gao2024magicdrivedit, wang2024drivedreamer, zhao2025drivedreamer2], we pursue an alternative: learning a 3D/4D world model as a high-fidelity, generative simulator [yang2024drivearena, yan2024drivingsphere]. At the heart of this paradigm lies the model’s ability to answer the critical question “What will happen if I take this action?”, by generating a holistic future scene that explicitly models the dynamic interactions between the ego-vehicle, the 3D/4D environment, and other agents.
While this generative approach holds immense promise, our work identifies a systemic and perilous flaw in these state-of-the-art models, which we term optimistic bias (Fig. 1(c)). When conditioned on an unsafe trajectory, these models refuse to predict the inevitable failure. Instead, they hallucinate an implausibly safe future. For instance, given a trajectory aimed at a pedestrian, the model will not forecast the collision; it might show the pedestrian vanishing entirely. Similarly, if commanded to drive onto a lawn, it may predict the grass magically turning into asphalt. This phenomenon reveals a fundamental breakdown in causal fidelity, where the model actively invents a safe outcome rather than forecasting the hazardous one commanded by the input. This makes it a dangerously unreliable critic for policy learning. We systematically quantify this bias with our proposed Risk Foreseeing Benchmark (RFB) (Sec. 4.2).
To this end, we introduce a plug-and-play closed-loop RL framework, AD-R1, to enhance the safety of any driving agent. At the heart of our framework is an internal critic, or more precisely, a high-fidelity simulation engine, powered by a specially designed impartial world model. This engine takes the agent’s planned trajectories as input, “dreams” of their future consequences in a high-fidelity simulation, and provides dense, safety-critical feedback. This allows the agent to learn from a vast spectrum of imagined mistakes in a safe, offline setting, effectively honing its policy to navigate away from potential failures.
The success of this framework, however, is predicated on a world model that can be honest about danger. Thus, our primary technical contribution is to build an Impartial World Model (IWM) that is trained to be veridical about the outcomes of both safe and unsafe actions alike. We build the IWM by tackling the optimistic bias at its source with Counterfactual Synthesis. This data pipeline treats real-world driving logs as editable stages, programmatically generating a rich curriculum of physically-consistent and causally-faithful ego-trajectories that lead to plausible failures. To further enforce causal fidelity, we introduce model-centric refinements, including a Trajectory-Aware Gating module and an Ego-Trajectory Fidelity Loss, which ensure the model’s predictions remain faithful to the commanded trajectory, even when it leads to a hazardous outcome.
This resulting IWM is then integrated as the core of our RL-based critic. Its ability to realistically simulate negative outcomes unlocks a new level of granularity and physical grounding in our reward modeling. Instead of relying on simple heuristics, we can now formulate precise, multi-faceted reward signals directly from the predicted 4D occupancy. This allows us to heavily penalize not only overt failures like collisions and off-road excursions but also subtle. unsafe behaviors such as violating spatio-temporal buffers or ignoring vertical clearance constraints, providing a rich and reliable signal for policy refinement.
The main contributions of our work are as follows:
  • We introduce AD-R1, a novel RL framework that unlocks policy refinement directly in the full 4D spatio-temporal domain. It leverages a generative occupancy world model as a high-fidelity ”dream engine,” enabling agents to learn safely from a vast spectrum of imagined failures—a critical capability for real-world autonomy.
  • To enable this, we pioneer the systematic diagnosis of the “Optimistic Bias” in world models. We forge the Risk Foreseeing Benchmark (RFB), the first benchmark designed to rigorously quantify a model’s ability to predict catastrophic failure, providing a critical new tool for the community.
  • Our solution is the Impartial World Model (IWM), engineered to be a veridical judge of risk. We build it via two synergistic thrusts: a data-centric pipeline, Counterfactual Synthesis, to teach it a rich curriculum of plausible failures; and model-centric refinements (TAG, Fidelity Loss) to ensure its predictions remain causally faithful to the agent’s commands.
  • Extensive experiments on navsim demonstrate AD-R1’s plug-and-play effectiveness across two baselines, boosting planning PDSM by 1.7% absolute.
2 Related Work
2.1 World Models for Autonomous Driving
The concept of a world model as a generative data engine to create high-fidelity, controllable driving videos has recently seen remarkable success [gao2023magicdrive, gao2024magicdrivedit, wang2024drivedreamer, zhao2025drivedreamer2]. Frameworks like DriveArena [yang2024drivearena] and DrivingSphere [yan2024drivingsphere] have pushed the boundaries of generative closed-loop simulation. Beyond general video synthesis, a parallel and crucial application of world models lies in their use as explicit, action-conditioned simulators for policy learning and planning. Approaches have explored everything from future video prediction [gao2024vista, jia2023adriver, zhang2025epona] to, more recently, explicit 3D/4D representations [kong20253d, yan2025olidm]. Voxel-based world models, in particular, have demonstrated remarkable capabilities in trajectory-conditioned future state prediction, capturing fine-grained 3D geometry and 4D spatio-temporal dynamics [bian2024dynamiccity, gu2024dome, shi2025come, zheng2024occworld, wang2024occsora, liao2025i2]. A parallel line of work leverages world models not as explicit simulators, but to enhance policy learning through auxiliary tasks, such as learning a latent scene representation [law, zheng2025world4drive] or predicting future BEV features [wote]. While this ensures fidelity for on-distribution scenarios, we argue that it inherently creates the ”optimistic bias”—a systemic failure to forecast the consequences of unsafe actions. This data-induced limitation makes them unreliable for safety-critical tasks like RL-based policy refinement. Our work is among the first to systematically diagnose and address this fundamental flaw, transforming the world model from a passive scene predictor into an active and impartial judge of safety by explicitly training it on a curriculum of synthesized failures.
2.2 End-to-end Autonomous Driving
End-to-end autonomous driving aims to learn a single unified policy from raw sensors to control. While modern architectures based on Transformers [hu2023uniad, jiang2023vad, chen2024vadv2, chitta2022transfuser, li2024hydra] and Diffusion Models [liao2025diffusiondrive, xing2025goalflow] have achieved significant success, they largely follow an Imitation Learning (IL) paradigm. A fundamental limitation of pure IL is its reliance on the expert data distribution; models struggle with causal confusion and are prone to failures in long-tail scenarios due to distributional shift. To overcome these challenges, a growing consensus is forming around the need to incorporate closed-loop Reinforcement Learning (RL). Early efforts explored model-free RL in simulators like CARLA [dosovitskiy2017carla, renz2024carllava], but sample inefficiency remains a major hurdle for real-world application. Consequently, model-based RL has become a more promising direction. Approaches like Think2Drive [li2024think2drive] and Raw2Drive [yang2025raw2drive] learn a world model within the CARLA simulator, but this paradigm inherits the simulator’s own sim-to-real gap. Some efforts used simulators with non-reactive 2D layouts [li2025recogdrive], their inability to model complex 3D/4D interactions is a major drawback. A significant breakthrough came with reconstruction-based simulation, such as RAD [gao2025rad], which leverages 3D Gaussian Splatting for photo-realism.
Instead of reconstructing a single, static scene, we utilize a learned world model to generate a vast diversity of potential future outcomes in response to the agent’s actions. Our primary contribution is an impartial world model that acts as this reliable generative simulator, or internal critic. This enables a generative world model-based RL framework where the agent can effectively and safely learn from imagined mistakes, leading to a demonstrably safer driving policy.
3 Methodology
Figure 2: Overview of AD-R1. Our framework involves two stages: (a) Impartial World Model Training: We first introduce a Counterfactual Synthesis pipeline that decomposes real scenes and programmatically generates unsafe trajectories to create Synthetic Negative Data. Our Impartial World Model is then trained on a mix of this synthetic failure data and real-world safe data. (b) Reinforcement Learning Post-training: The trained model acts as an internal critic. It takes candidate trajectories from a pre-trained agent, ”dreams” of the 4D future outcomes, and our 4D Rewarded Modeling module computes a dense reward based on the imagined world. This provides a strong Policy Loss to refine the agent’s safety and robustness through imagined failures.
3.1 Overview and Problem Formulation
Our goal is to refine a pre-trained end-to-end driving policy, $\pi$, which maps sensor observations $\mathcal{O}_{t}$ and a command $\mathcal{G}$ to a future ego-trajectory $\mathcal{T}_{\text{ego}}$. While often trained with imitation learning, such policies struggle in long-tail events [li2025recogdrive, gao2025rad], we formulate this as a reinforcement learning (RL) problem, where the policy $\pi$ is optimized to maximize the expected future reward from a reward function $R$:
$\pi^{*}=\arg\max_{\pi}\mathbb{E}_{\mathcal{T}_{\text{ego}}\sim\pi(\mathcal{O}_{t},\mathcal{G})}\left[R(\mathcal{T}_{\text{ego}},\mathcal{O}_{t})\right],$ (1)
where the reward function $R$ evaluates the quality of a proposed trajectory $\mathcal{T}_{\text{ego}}$.
To obtain this reward signal safely and efficiently, existing methods [dauner2024navsim, li2025recogdrive, gao2025rad] rely on external simulators. However, as discussed, these are limited by sim-to-real gaps in graphics engines, the non-interactive nature of 2D layout simulators, or the static, costly nature of reconstruction-based environments. In a departure from these paradigms, we introduce a learned world model, $\mathcal{M}$, to serve as a high-fidelity, internal generative simulator. $\mathcal{M}$ takes the historical observations $\mathcal{O}_{\leq t}$ and a candidate ego-trajectory $\mathcal{T}_{\text{ego}}$ as input, and predicts the corresponding future evolution of the scene, typically represented as a sequence of future 3D occupancy grids $\{\hat{\mathcal{O}}_{t+k}\}_{k=1}^{K}$:
$\{\hat{\mathcal{O}}_{t+k}\}_{k=1}^{K}=\mathcal{M}(\mathcal{O}_{\leq t},\mathcal{T}_{\text{ego}}).$ (2)
The reward $R$ is then computed directly from this predicted future. Therefore, the entire framework hinges on the model’s fidelity. A good world model must faithfully predict the consequences of any given trajectory, especially unsafe ones, to provide a reliable reward signal.
As shown in Fig. 2, our method is thus composed of two key components: (1) the design of a novel Impartial World Model (detailed in Sec. 3.2), and (2) its integration into our RL framework for policy refinement (detailed in Sec. 3.3).
3.2 Impartial Occupancy World Model
While state-of-the-art world models exhibit impressive forecasting performance, they suffer from the “optimistic bias” discussed previously. To address this, we build an occupancy-based prediction model designed to capture the complex 3D geometry and 4D spatio-temporal dynamics of the driving scene. Crucially, this model is fundamentally engineered to overcome this bias. We first outline the backbone architecture before detailing our synergistic data-centric and model-centric refinements.
3.2.1 Backbone Architecture.
Our impartial world model builds upon a powerful and efficient architecture inspired by I²-World [liao2025i2world]. The model operates in a latent space for computational efficiency, following a two-stage process:
  • Scene Tokenization: An encoder, termed the I²-Scene Tokenizer, first compresses the sequence of high-dimensional historical occupancy grids $\mathcal{O}_{\leq t}$ into a compact sequence of latent tokens $\mathcal{B}_{\leq t}$. This tokenizer is designed to efficiently capture both the static spatial details within each frame (intra-scene) and the dynamic temporal evolution across frames (inter-scene).
  • Autoregressive Forecasting: A transformer-based forecaster, the I²-Former, then autoregressively predicts a sequence of future latent tokens $\{\hat{\mathcal{B}}_{t+k}\}_{k=1}^{K}$. This forecaster consists of an encoder that processes the historical tokens and the candidate ego-trajectory, and a decoder that generates the future tokens. The final sequence of occupancy grids $\{\hat{\mathcal{O}}_{t+k}\}_{k=1}^{K}$ is produced by decoding these latent predictions.
This robust two-stage architecture provides the foundation upon which we introduce our key refinements to instill the vital capability of imagining failure.
3.2.2 Counterfactual Data Synthesis.
Existing real-world datasets [caesar2020nuscenes, sun2020waymo], which form the foundation of most world models, primarily consist of professional trajectories recorded in safe driving scenarios. The inherent lack of failure examples in this data is the primary cause of optimistic bias. We address this at its source by proposing a systematic data synthesis pipeline to create a diverse and structured curriculum of counterfactual negative samples.
Our pipeline treats each safe scenario from the dataset as an editable stage. Given the historical occupancy sequence $O_{\leq t}$, we first decompose the current frame $O_{t}$. This decomposition leverages the ground-truth 3D annotations provided by the nuScenes dataset [caesar2020nuscenes, tian2023occ3d] to isolate the static environment, dynamic agents, and the original ego-vehicle’s state using clustering [schubert2017dbscan]. This provides the necessary context to generate plausible, scene-aware failure trajectories. We acknowledge that this reliance on annotations is a current limitation, though we believe it could be replaced by a high-fidelity perception system in the future.
Next, we programmatically generate a new, unsafe ego-trajectory $\tilde{T}_{\text{ego}}=\{\mathbf{p}_{t+k}\}_{k=1}^{K}$. All unsafe trajectories are generated by a unified kinematic model. This model steers the ego-vehicle from its current state toward a designated target point $\mathbf{p}_{\text{target}}$, which is selected based on the desired failure mode. The future position $\mathbf{p}_{t+k}$ at each timestep $k$ is determined by blending the vehicle’s historical momentum with a vector towards the target:
$\mathbf{v}_{\text{target}}=\frac{\mathbf{p}_{\text{target}}-\mathbf{p}_{t}}{\|\mathbf{p}_{\text{target}}-\mathbf{p}_{t}\|}$ (3)
$\mathbf{p}_{t+k}=\mathbf{p}_{t}+\alpha_{k}\cdot\Delta d_{k}\cdot\mathbf{h}_{\text{hist}}+\beta_{k}\cdot\Delta d_{k}\cdot\mathbf{v}_{\text{target}}+\mathcal{N}(0,\sigma^{2})$ (4)
where $\mathbf{h}_{\text{hist}}$ is the historical heading, $\mathbf{v}_{\text{target}}$ is the unit vector towards the target, $\Delta d_{k}$ is the step distance, and $\alpha_{k},\beta_{k}$ are time-varying blending coefficients. These coefficients are designed to create a smooth, plausible maneuver that starts from the current trajectory and progressively deviates towards the target, mimicking a realistic driving action.
We construct a hierarchical curriculum of failures by defining strategies for selecting the target point $\mathbf{p}_{\text{target}}$:
Ego-State Failure (Off-Road Events). To simulate leaving the roadway, $\mathbf{p}_{\text{target}}$ is selected as a point just outside the nearest drivable area boundary. Applying our kinematic model (Eq. 4) generates a smooth off-road drift.
Interaction with Static Environment (Penetration). The next level of complexity involves interacting with non-movable parts of the world. We generate trajectories that lead to the ego-vehicle penetrating static obstacles such as buildings, barriers, or vegetation. This is achieved by planning a path directly into a nearby static object identified from the decomposed scene map. This teaches the model the concept of solid objects and physical impassability.
Interaction with Dynamic Agents (Collisions). To simulate a dynamic collision with another agent, $\mathbf{p}_{\text{target}}$ is set to a future predicted position. For simplicity, we assume a linear or constant-velocity prediction for the target agent, and our kinematic model generates an intercept trajectory.
By using a single, flexible trajectory generation model applied to a structured hierarchy of targets, we can efficiently synthesize a rich and diverse curriculum. Finally, we synthesize the new ground-truth future occupancy sequence $\{\mathcal{O}^{\prime}_{t+k}\}_{k=1}^{K}$. For each future timestep $k$, we compute the rigid-body transformation $\mathbf{T}_{k}$ that maps voxel coordinates from the new counterfactual ego-frame back to the original one. The new future grid $\mathcal{O}^{\prime}_{t+k}$ is then generated by applying this transformation to the original grid $\mathcal{O}_{t+k}$, using interpolation for non-integer coordinates:
$\mathcal{O}^{\prime}_{t+k}(\mathbf{v}^{\prime})=\mathcal{O}_{t+k}(\mathbf{T}_{k}(\mathbf{v}^{\prime})).$ (5)
This operation, using interpolation for non-integer coordinates, ensures that all complex spatial relationships are preserved with high physical fidelity. By training on this rich curriculum of synthesized negative samples, we enable our world model $\mathcal{M}$ to learn the crucial causal link between unsafe ego-trajectories and their corresponding failure outcomes, such as collisions and off-road events.
3.2.3 Model-Centric Refinements
Trajectory-Aware Gating (TAG). To further enhance the model’s fidelity to control input, we introduce a lightweight, plug-and-play gating module. Specifically, we first rasterize the input ego-trajectory $\mathcal{T}_{\text{ego}}$ into a 2D spatial mask, which a small CNN processes into a gating map, $\mathcal{A}_{\text{traj}}\in[0,1]^{h\times w}$. This map dynamically modulates the intermediate scene features $\mathcal{F}_{\text{scene}}$ via element-wise multiplication:
$\mathcal{F}_{\text{gated}}=\mathcal{F}_{\text{scene}}\odot\mathcal{A}_{\text{traj}}.$ (6)
This forces the model to prioritize features along the commanded path, preventing it from “ignoring” an obstacle to invent a safe outcome.
Ego-Trajectory Fidelity Loss. The final piece is a novel loss function that penalizes any deviation of the predicted ego-vehicle from its commanded trajectory. For each predicted future timestep $k$, we decode the predicted occupancy map $\hat{\mathcal{O}}_{t+k}$ and compute the geometric center of all voxels classified as the ego-vehicle, denoted as $\hat{p}_{\text{ego}}^{t+k}$. We then introduce an Ego-Trajectory Fidelity Loss, $\mathcal{L}_{\text{fidelity}}$, which measures the L2 distance between this predicted center and the ground-truth position from the input control trajectory, $p_{\text{ego}}^{t+k}$:
$\mathcal{L}_{\text{fidelity}}=\sum_{k=1}^{K}w_{k}\left\|\hat{p}_{\text{ego}}^{t+k}-p_{\text{ego}}^{t+k}\right\|_{2}^{2}.$ (7)
This loss acts as a strong regularizer, directly discouraging the model from hallucinating a future where the ego-vehicle autonomously “swerves” to avoid a collision. It enforces the principle that the world model’s role is to predict, not to act.
Final Training Objective. Our impartial world model is trained on a mixed dataset of real and synthesized counterfactual data by augmenting the original baseline objective $\mathcal{L}_{\text{baseline}}$ [liao2025i2] with our proposed fidelity loss:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{baseline}}+\lambda\mathcal{L}_{\text{fidelity}},$ (8)
where $\lambda$ is a balancing hyperparameter. The model is trained on a mixed dataset comprising both original safe scenarios and our synthesized failure scenarios.
3.3 RL with Future Imagination
3.3.1 Training Paradigm
Our training paradigm leverages our impartial world model as a high-fidelity simulator to refine a pre-trained policy $\pi$ in a closed-loop, yet entirely offline, manner. To enable stable and efficient exploration beyond simple imitation, we introduce a Group Relative Policy Optimization (GRPO) algorithm [guo2025deepseek], adapted for trajectory refinement.
Concretely, for a given observation $\mathcal{O}_{t}$, we refine a diffusion-based driving policy $\pi_{\theta}$. Following [li2025recogdrive], we formulate the diffusion denoising process as a finite-horizon Markov Decision Process (MDP), where the diffusion planner acts as a stochastic policy. Starting from Gaussian noise $\mathcal{T}_{T}\sim\mathcal{N}(0,\mathbf{I})$, the policy gradually denoises over $T$ steps to generate a full trajectory $\mathcal{T}_{\text{ego}}=\mathcal{T}_{0}$. To enable exploration beyond imitation, we sample a group of $G$ trajectory rollouts, $\{\mathcal{T}_{\text{ego}}^{(i)}\}_{i=1}^{G}$. Each trajectory is then passed to our world model $\mathcal{M}$ to “dream” of its consequences, yielding a corresponding future occupancy sequence $\hat{S}^{(i)}=\mathcal{M}(\mathcal{O}_{t},\mathcal{T}_{\text{ego}}^{(i)})$. We then compute a scalar reward $r_{i}=\mathcal{R}(\hat{S}^{(i)})$ (Sec. 3.3.2) for each imagined rollout using our reward function.
Unlike prior approaches that rely on simple surrogate rewards, our physically-grounded rewards provide realistic feedback on safety and comfort. We treat the entire group of rollouts as a single-step decision process. To obtain a stable learning signal, we compute a group-standardized advantage for each trajectory:
$\hat{A}_{i}=\frac{r_{i}-\text{mean}(\{r_{j}\}_{j=1}^{G})}{\text{std}(\{r_{j}\}_{j=1}^{G})},\quad i=1,\dots,G.$ (9)
This normalization centers the rewards and scales them by their variance, focusing the policy update on the relative quality of trajectories within the sampled group.
Finally, we compute the policy loss. The objective is composed of two main terms: a reinforcement learning loss that encourages reward-seeking behavior, and a behavior cloning loss that regularizes the policy to prevent catastrophic forgetting.
The RL loss, based on the GRPO formulation, is computed over the group of sampled trajectories:
$\mathcal{L}_{\text{RL}}=-\frac{1}{G}\sum_{i=1}^{G}\log\pi_{\theta}(\mathcal{T}_{\text{ego}}^{(i)}\mid\mathcal{O}_{t})\hat{A}_{i}.$ (10)
This term updates the policy to increase the likelihood of trajectories that yield a higher relative advantage within the group.
To prevent the policy from deviating too far from the sensible priors of the pre-trained model, we add a behavior cloning (BC) loss. This term anchors the refined policy $\pi_{\theta}$ to the original reference policy $\pi_{\text{ref}}$:
$\mathcal{L}_{\text{BC}}=\mathbb{E}_{\mathcal{T}_{\text{ref}}\sim\pi_{\text{ref}}}\left[-\log\pi_{\theta}(\mathcal{T}_{\text{ref}}\mid\mathcal{O}_{t})\right].$ (11)
The final training objective is a weighted sum of these two components:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{RL}}+\lambda\mathcal{L}_{\text{BC}},$ (12)
where $\lambda$ is a hyperparameter that controls the strength of the regularization. Through this GRPO-based objective, the driving policy learns to generate safer and more comfortable trajectories in a purely offline setting, effectively learning from a vast spectrum of imagined failures.
3.3.2 4D Reward Modeling
[…]
3.3.3 Safety-Critical Penalties
The foremost priority is to penalize any future that our model predicts as unsafe.
Volumetric Collision Penalty ($\boldsymbol{\mathcal{R}_{\text{coll}}}$): A collision is the most severe failure event. We go beyond a simple binary penalty by quantifying the severity of the predicted collision. For any predicted timestep $t$, we identify the set of voxels occupied by the ego vehicle, $\mathcal{V}_{\text{ego}}$, and any other dynamic agent $k$ (e.g., another vehicle or a pedestrian), $\mathcal{V}_{\text{agent},k}$. A collision occurs if their intersection is non-empty. We define the Volumetric Collision IoU (VC-IoU) to quantify the spatial overlap:
$\text{VC-IoU}_{k}=\frac{|\mathcal{V}_{\text{ego}}\cap\mathcal{V}_{\text{agent},k}|}{|\mathcal{V}_{\text{ego}}\cup\mathcal{V}_{\text{agent},k}|}.$ (13)
This metric naturally captures the extent of the overlap, providing a proxy for impact severity. The collision penalty then heavily penalizes any contact while scaling with this severity:
$\mathcal{R}_{\text{coll}}=-\sum_{t,k}\left[C_{\text{base}}\cdot\mathbb{I}(\text{VC-IoU}_{k}>0)+C_{\text{iou}}\cdot\text{VC-IoU}_{k}\right],$ (14)
where $\mathbb{I}(\cdot)$ is the indicator function, $C_{\text{base}}$ is a large constant penalty for any contact, and $C_{\text{iou}}$ weights the severity.
Off-Road and Sidewalk Penalty ($\boldsymbol{\mathcal{R}_{\text{offroad}}}$): Leveraging semantic predictions, we penalize the ego vehicle for occupying non-drivable areas. We identify the set of voxels under the ego vehicle’s footprint, $\mathcal{V}_{\text{footprint}}$. The penalty is proportional to the number of these voxels classified with non-permissible semantics:
$\mathcal{R}_{\text{offroad}}=-\sum_{t}|\{\mathbf{v}\in\mathcal{V}_{\text{footprint}}\mid\texttt{Sidewalk}$
$\in\mathcal{O}_{t}(\mathbf{v})\lor\texttt{Vegetation}\in\mathcal{O}_{t}(\mathbf{v})\}|.$ (15)
3.3.4 3D/4D-Aware Comfort & Rule-Adherence
Here, we introduce the design of rewards uniquely enabled by the 3D and temporal awareness of our model, showcasing clear advantages over 2D BEV.
Vertical Clearance Penalty ($\boldsymbol{\mathcal{R}_{\text{clearance}}}$): Safe driving requires awareness of vertical space, to which 2D BEV representations are blind. Our 3D model prevents collisions with low-hanging structures (e.g., bridges, tunnels). We define a 3D “safety bubble” above the ego vehicle, $\mathcal{V}_{\text{bubble}}$, and penalize its intersection with static obstacles:
$\mathcal{R}_{\text{clearance}}=-\sum_{t}\left|\{\mathbf{v}\in\mathcal{V}_{\text{bubble}}\mid\texttt{Building}\in\mathcal{O}_{t}(\mathbf{v})\}\right|.$ (16)
This provides a crucial safety constraint impossible to model from a BEV perspective alone.
Drivable Surface Stability Reward ($\boldsymbol{\mathcal{R}_{\text{stability}}}$): A region labeled “drivable” in 2D may be a steep curb or uneven road in 3D. We reward the agent for choosing stable, flat surfaces by penalizing high variance in the Z-coordinate (height) of the drivable surface voxels directly beneath the ego vehicle’s footprint:
$\mathcal{R}_{\text{stability}}=$ $-\sum_{t}\text{Var}\left(Z(\mathbf{v})\right)\qquad\text{for }$
$\mathbf{v}\in\{\text{voxels below }$ $\mathcal{V}_{\text{footprint}}\mid\texttt{Drivable}\in\mathcal{O}_{t}(\mathbf{v})\}.$ (17)
This encourages a smoother and safer ride.
3.3.5 Task-Oriented Rewards
Finally, to ensure the agent makes progress, we include standard task-oriented rewards.
Progress Reward ($\boldsymbol{\mathcal{R}_{\text{progress}}}$): Reward for advancing along a pre-defined route or towards a goal.
Velocity Reward ($\boldsymbol{\mathcal{R}_{\text{velocity}}}$): Reward for maintaining a target speed, typically the local speed limit.
The final reward passed to the RL algorithm is the weighted sum of all components:
$\mathcal{R}_{\text{total}}=w_{\text{coll}}\mathcal{R}_{\text{coll}}+w_{\text{offroad}}\mathcal{R}_{\text{offroad}}+w_{\text{clear}}\mathcal{R}_{\text{clearance}}$
$+\dots+w_{\text{prog}}\mathcal{R}_{\text{progress}}.$ (18)
By combining these multi-faceted reward signals, our framework fully exploits the rich, predictive power of the impartial 4D world model, enabling the learning of a driving policy that is not only effective but also demonstrably safer and more physically-grounded.
4 Experiments
4.1 Experimental Setup
Our impartial world model is trained on a mixed dataset. The primary source of safe driving scenarios is the nuScenes dataset [caesar2020nuscenes], providing rich sensor data and annotations. This is augmented with failure scenarios generated via our Counterfactual Synthesis pipeline. We use a data split of 80% real nuScenes data and 20% our synthesized counterfactual data. This ratio was determined through empirical validation to provide a sufficient number of failure examples without significantly shifting the model’s prior away from realistic, safe driving dynamics. For the closed-loop evaluation of the final driving policy, we use the official NavSim benchmark [dauner2024navsim], which offers a standardized environment. We evaluate our method at two levels: the fidelity of the world model and the performance of the refined driving policy. Detailed Metric and Implementation are included in the supplementary.
4.2 Quantitative Analysis
AD-R1 Excels at Foreseeing Risk. As shown in Tab. 1, standard world models struggle significantly on our RFB benchmark. DOME [gu2024dome] and $I^{2}$-world [liao2025i2], trained only on safe data, exhibit a strong optimistic bias, resulting in very low $f$-IoU and DAF. In contrast, our IWM, trained with our synthetic counterfactual data and model-centric refinements, achieves dramatically higher scores. This quantitatively validates our primary claim: our impartial world model can faithfully “imagine failure” where other models hallucinate safety.
Table 1: World Model Performance on the Risk Foreseeing Benchmark (RFB). AD-R1 significantly outperforms the baseline in predicting the outcomes of unsafe trajectories, demonstrating its ability to overcome optimistic bias.
World Model G-IoU (%) $\uparrow$ f-IoU $\uparrow$ DAF $\uparrow$
DOME [gu2024dome] 21.01 14.21 8.75
I²-World [liao2025i2world] 30.91 21.44 14.21
IWM (Ours) 40.21 45.91 25.10
Table 2: Closed-Loop Performance on NavSim. Our post-training refinement framework, powered by AD-R1, significantly improves the safety and planning quality of two state-of-the-art baseline agents.
Method Scheme WM NC $\uparrow$ DAC $\uparrow$ TTC $\uparrow$ Comf. $\uparrow$ EP $\uparrow$ PDMS $\uparrow$
Constant Velocity 68.0 57.8 50.0 100 19.4 20.6
Ego Status MLP 93.0 77.3 83.6 100 62.8 65.6
VADv2 [chen2024vadv2] IL 97.2 89.1 91.6 100 76.0 80.9
DrivingGPT [chen2025drivinggpt] IL 98.9 90.7 94.9 100 79.7 82.4
UniAD [hu2023uniad] IL 97.8 91.9 92.9 100 78.8 83.4
TransFuser [chitta2022transfuser] IL 97.7 92.8 92.8 100 79.2 84.0
Hydra-MDP [li2024hydra] IL 98.3 96.0 94.6 100 78.7 86.5
LAW [law] IL+SSL 97.2 93.3 91.9 100 78.8 83.8
World4Drive [zheng2025world4drive] IL 97.4 94.3 92.8 100 79.9 85.1
Trajhf [li2025trajhf] IL+RL 96.6 96.6 92.1 100 84.5 87.6
WoTE [wote] IL 98.5 96.8 94.9 99.9 81.9 88.3
DriveDPO [shang2025drivedpo] IL+RL 98.5 98.1 94.8 99.9 84.3 90.0
GoalFlow [xing2025goalflow] IL 98.4 98.3 94.6 100 85.0 90.3
Recogdrive [li2025recogdrive] IL+RL 97.9 97.3 94.9 100 87.3 90.8
DiffusionDrive [liao2025diffusiondrive] IL 98.2 96.2 94.7 100 82.2 88.1
+ AD-R1 (Ours) RL 98.5 97.5 94.6 100 84.8 89.8 (+1.7)
Recogdrive-stage2 [li2025recogdrive] IL 98.1 94.7 94.2 100 80.9 86.5
+ AD-R1 (Ours) RL 98.74 97.8 94.8 100 87.5 91.9 (+1.1)
Post-Training Refinement Boosts Safety and Performance. Tab. 2 presents the core results of our paper. Applying our plug-and-play refinement framework yields substantial improvements across both baselines. For both diffusion-policy based method, DiffusionDrive [liao2025diffusiondrive] and VLA-based method Recogdrive [li2025recogdrive], we observe a stable improvement of PDMS scores in closed-loop simulations. This demonstrates that learning from imagined failures directly translates to safer real-world behavior. Furthermore, this enhanced safety does not come at the cost of performance. Our method improves the PDSM score for both agents, with an absolute boost of 1.7% and 1.1%, respectively.
4.3 Ablation Studies
Effect of Core Components. We conduct a comprehensive ablation study to dissect the contribution of each of our proposed components, with results presented in Tab. 3. We begin with a strong IL-trained baseline, DiffusionDrive. Our first key finding is that naively applying a standard, optimistic world model (e.g., I²-World) for RL-based refinement is actively harmful, causing a significant drop in the PDMS score (88.1 $\rightarrow$ 85.3).
In contrast, incrementally adding our contributions demonstrates a clear and consistent path to improvement. Introducing Counterfactual Synthesis (CS) provides the most substantial boost, lifting the PDMS score to 89.1. This highlights that teaching the model to recognize failure is the single most critical factor. Subsequently, adding our Model-centric Refinements (MR) (TAG and Fidelity Loss) further improves the score to 89.3 by enforcing stricter causal fidelity. Finally, leveraging our physically-grounded 4D Rewards (R) in the refinement loop yields the best overall performance (89.8 PDMS). This validates that each of our contributions—curing the optimistic bias, enforcing causal fidelity, and utilizing detailed 3D/4D rewards—plays a crucial and synergistic role in successfully and safely refining the driving policy.
Table 3: Ablation study on the components of AD-R1 and our RL framework. Results are reported on the G-IoU of RFB (for the world model) and NavSim PDMS (for the final policy).
Configuration G-IoU $\uparrow$ PDMS $\uparrow$
Baseline (DiffusionDrive) - 88.1
+ $i^{2}$-world 30.91 85.3
+ CS 37.22 89.1
+ CS + MR 40.21 89.3
+ CS + MR + R 40.21 89.8
4.4 Qualitative Analysis
Visualizing Failure Imagination. Fig. 4 provides a stark qualitative comparison. We feed an identical unsafe off-road trajectory into a standard world model [gu2024dome] and into AD-R1. The standard model hallucinates a safe future, with the disappearing of the greenbelt in the middle of the road. In contrast, IWM faithfully predicts the off-road scene, generating an imagined future that provides a clear and correct penalty signal. More visualizations are incuded in the supplementary.
Visualizing Policy Refinement In Figure 3, we show the behavior of a baseline agent versus our refined agent in a critical scenario.
Figure 3: Behavior of an agent with and without AD-R1 refinement. Left: The original agent’s plan results in a collision or off-road. Right: Our refined agent safely avoids the hazard.
Figure 4: Qualitative comparison of world models.Top: The reference synthetic data. Middle: DOME[gu2024dome]’s optimistic hallucination. Bottom: AD-R1’s faithful off-road prediction.
5 Conclusion
Reinforcement learning for driving is stalled by a fundamental flaw: the “optimistic bias” of world models that cannot predict failure. We solve this by introducing Counterfactual Synthesis to train an Impartial World Model that faithfully predicts the consequences of unsafe actions. As an internal critic in our AD-R1 framework, this model enables agents to learn from imagined mistakes, drastically reducing safety violations in closed-loop simulations. We demonstrate that teaching a model to dream of danger is not a limitation, but a prerequisite for building truly safe and intelligent autonomous agents. This work demonstrates that the path to robust autonomy is not to shield agents from failure, but to grant them the foresight to imagine it.

### v2

@@
 2.1 World Models for Autonomous Driving
-The concept of a world model as a generative data engine to create high-fidelity, controllable driving videos has recently seen remarkable success [gao2023magicdrive, gao2024magicdrivedit, wang2024drivedreamer, zhao2025drivedreamer2].
+The concept of a world model as a generative data engine to create high-fidelity, controllable driving videos has recently seen remarkable success [gao2023magicdrive, gao2024magicdrivedit, wang2024drivedreamer, zhao2025drivedreamer2, yan2025rlgf, liang2025worldlens, tang2025omnigen].
 Frameworks like DriveArena [yang2024drivearena] and DrivingSphere [yan2024drivingsphere] have pushed the boundaries of generative closed-loop simulation.
@@
 Approaches have explored everything from future video prediction [gao2024vista, jia2023adriver, zhang2025epona] to, more recently, explicit 3D/4D representations [kong20253d, yan2025olidm].
-Voxel-based world models, in particular, have demonstrated remarkable capabilities in trajectory-conditioned future state prediction, capturing fine-grained 3D geometry and 4D spatio-temporal dynamics [bian2024dynamiccity, gu2024dome, shi2025come, zheng2024occworld, wang2024occsora, liao2025i2].
+Voxel-based world models, in particular, have demonstrated remarkable capabilities in trajectory-conditioned future state prediction, capturing fine-grained 3D geometry and 4D spatio-temporal dynamics [bian2024dynamiccity, gu2024dome, shi2025come, zheng2024occworld, wang2024occsora, liao2025i2, deng2025best3dscenerepresentation, deng2025mcnslammultiagentcollaborativeneural, yan2022plug].
 A parallel line of work leverages world models not as explicit simulators, but to enhance policy learning through auxiliary tasks, such as learning a latent scene representation [law, zheng2025world4drive] or predicting future BEV features [wote].
@@
 This work demonstrates that the path to robust autonomy is not to shield agents from failure, but to grant them the foresight to imagine it.
+[…]
+Supplementary Material
+AD-R1: Closed-Loop Reinforcement Learning for End-to-End Autonomous Driving with Impartial World Models
+Supplementary Material
+Overview
+This supplementary material provides additional details to support the claims and ensure the reproducibility of the main paper. The content is organized as follows:
+  • Sec. A: Detailed network architecture, hyperparameters for the Impartial World Model (IWM), and the RL training configuration.
+  • Sec. B: Elaboration on the nuScenes-CF dataset and the mathematical formulation of Counterfactual Synthesis.
+  • Sec. C: Strict mathematical definitions for the Risk Foreseeing Benchmark (RFB).
+  • Sec. D: Additional quantitative results focusing on the sensitivity analysis of synthetic data ratios.
+  • Sec. E: Qualitative analysis and description of the attached demo.
+A Implementation Details
+A.1 Impartial World Model (IWM) Architecture
+Our IWM consists of a VQ-VAE-based tokenizer and a Transformer-based autoregressive forecaster based on I2-World [liao2025i2].
+4D Scene Tokenizer.
+We voxelize the local scene into a grid of spatial resolution $200\times 200$ and height resolution $16$.
+The physical range covered is $40\text{m},40\text{m}]$ in $X,Y$ and $1\text{m},5.4\text{m}]$ in $Z$, resulting in a voxel size of $0.4\text{m}\times 0.4\text{m}\times 0.4\text{m}$, following Occ-3D [tian2023occ3d].
+  • Encoder/Decoder: We utilize a 3D-VAE architecture with 4 downsampling stages.
+  • Vector Quantization: We use a codebook of size $N=512$ with an embedding dimension of $D=128$.
+4D Forecaster. The forecaster is a encoder-decoder Transformer tailored for 4D occupancy generation.
+  • Structure: 24 Attention Layers, 16 Attention Heads, Hidden Dimension $d_{model}=1024$.
+  • Context: The model takes 2 seconds of historical context (4 frames at 2Hz) and predicts 3 seconds into the future (6 frames).
+Training Details.
+We use AdamW optimizer with $1\times 10^{-3}$ as the base learning rate, 256 as the batch size.
+The Tokenizer is trained for 30 epochs while the Forecaster is trained for 50 epochs.
+All training is performed on $8\timesNVIDIA$ H20 GPUs.
+A.2 RL Training Configuration (AD-R1)
+We employ the Group Relative Policy Optimization (GRPO) algorithm adapted for continuous trajectory refinement.
+Training Setup.
+  • Policy Network: We fine-tune the denoising head of the pre-trained DiffusionDrive [liao2025diffusiondrive] and ReCogDrive [li2025recogdrive].
+  • Group Sampling ($G$): Due to the computational cost of 4D world model rollouts, we set the group size $G=64$ and $G=8$ for DiffusionDrive and ReCogDrive, respectively.
+This provides a sufficient baseline variance for advantage estimation while fitting within GPU memory constraints.
+  • Optimization: We use the AdamW optimizer with a learning rate of $1\times 10^{-5}$ and $4\times 10^{-5}$ for DiffusionDrive and ReCogDrive.
+  • Hardware: Training is performed on 8 $\times$ NVIDIA H20 GPUs. The refinement stage takes approximately 24 hours for 10 epochs on the navsim [dauner2024navsim] training set.
+A.3 Reward Function Details
+The reward shaping is critical for guiding the agent towards safety without compromising progress. Table 4 details the specific weights used in Eq. (18) of the main paper.
+B The nuScenes-CF Dataset
+B.1 Kinematic Trajectory Generation
+To generate the unsafe ego-trajectory $\tilde{\mathcal{T}}_{ego}$ for Counterfactual Synthesis, we employ a Blending Kinematic Model.
+Let $\mathbf{p}_{t},\mathbf{v}_{t}$ be the vehicle’s current state and $\mathbf{p}_{target}$ be the designated collision point (e.g., the center of another vehicle or a wall).
+The synthesized position $\mathbf{p}_{t+k}$ at future timestep $k$ is computed as:
+$\mathbf{p}_{t+k}=(1-\gamma_{k})\cdot\mathbf{p}_{inertial}^{(k)}+\gamma_{k}\cdot\mathbf{p}_{intercept}^{(k)}$ (19)
+where:
+  • $\mathbf{p}_{inertial}^{(k)}$ represents the position if the vehicle continued with its historical momentum.
+  • $\mathbf{p}_{intercept}^{(k)}$ is the position required to linearly intercept the target $\mathbf{p}_{target}$.
+  • $\gamma_{k}\in[0,1]$ is a time-varying blending coefficient, defined as $\gamma_{k}=(k/K)^{2}$.
+The quadratic growth ensures a smooth deviation that mimics a loss of control or a gradual drift, rather than an unrealistic sharp turn.
+B.2 Dataset Statistics and Sampling Strategy
+We constructed a dedicated pool of counterfactual scenarios consisting of 1,800 synthetic clips derived from the nuScenes training set.
+For evaluation, we curated a separate set of 450 clips to serve as the Risk Foreseeing Benchmark (RFB).
+Training Sampling Strategy.
+It is important to note that the ”80:20 ratio” refers to the data distribution seen by the model during training, not the absolute size of the datasets.
+We employ a weighted sampling strategy: in each training iteration, 80% of the samples are drawn from the original real-world nuScenes dataset (safe expert data), and 20% are sampled from our synthetic counterfactual pool (unsafe failure data).
+This ensures the model maintains strong priors on realistic scene dynamics while receiving sufficient exposure to rare failure modes.
+The distribution of failure types within the synthetic pool is balanced as follows:
+  • Dynamic Collisions (40%): Inter-vehicle collisions and vehicle-pedestrian accidents.
+  • Static Collisions (30%): Collisions with barriers, traffic cones, and buildings.
+  • Off-Road/Sidewalk (30%): Excursions into non-drivable zones.
+B.3 Counterfactual Example
+We visualize examples about synthetic counterfactual occupancy data to show Penetration (Fig. 6), Off-road (Fig. 5) and Collision (Fig. 7).
+Figure 5: Examples of synthetic unsafe trajectories for the Counterfactual Data (Off-road). Red dot indicates the end of the trajectory while the green one indicates the start point.
+Figure 6: Examples of synthetic unsafe trajectories for the Counterfactual Data (Penetration). Red dot indicates the end of the trajectory while the green one indicates the start point.
+Figure 7: Examples of synthetic unsafe trajectories for the Counterfactual Data (Collision). Red dot indicates the end of the trajectory while the green one indicates the start point.
+C Experimental Setup and Benchmarks
+C.1 Risk Foreseeing Benchmark (RFB)
+To quantify the core contribution of our AD-R1 for overcoming the optimistic bias, we evaluate world models on our proposed Risk Foreseeing Benchmark (RFB).
+The RFB consists of a curated set of 450 challenging scenarios (clips) with ground-truth unsafe ego-trajectories spanning 3 seconds.
+Unlike standard benchmarks that evaluate on safe expert driving, RFB strictly tests the model’s ability to predict negative outcomes.
+We measure performance using the following three key metrics:
+  1. 1.
+    Global Scene Fidelity (G-IoU): This metric calculates the mean Intersection-over-Union (mIoU) across the entire scene (including static background and dynamic agents).
+It serves as a sanity check to penalize unrelated hallucinations and ensure the overall scene structure remains consistent.
+    $\text{{G}-IoU}=\frac{1}{|C|}\sum_{c\in C}\frac{|\hat{O}_{c}\cap O_{c}|}{|\hat{O}_{c}\cup O_{c}|}$ (20)
+    where $C$ represents the semantic classes in the scene.
+  2. 2.
+    Failure IoU (f-IoU): To strictly measure the prediction of danger, we define f-IoU.
+This metric focuses on the critical failure regions (e.g., collision zones surrounding the ego car).
+We define a local volume $V_{crit}$ centered around the ego-vehicle’s future trajectory.
+f-IoU calculates the IoU of occupied voxels specifically within this volume:
+    $\text{{f}-IoU}=\frac{|\hat{O}_{occ}\cap O_{occ}\cap V_{crit}|}{|\hat{O}_{occ}\cup O_{occ}\cap V_{crit}|}$ (21)
+    A high f-IoU indicates the model correctly predicts that ”something is there” (e.g., a wall or another car) directly in the ego-vehicle’s path, rather than hallucinating empty space.
+  3. 3.
+    Dynamic Agent Fidelity (DAF): Previous optimistic models tend to make other agents vanish (”ghosting”) to avoid collisions.
+DAF measures the average Instance-IoU of other dynamic agents (Vehicles, Pedestrians) to ensure that risk-inducing actors are preserved in the prediction.
+    $\text{DAF}=\frac{1}{N_{obj}}\sum_{i=1}^{N_{obj}}\text{IoU}(\hat{M}_{i},M_{i})$ (22)
+    where $M_{i}$ is the occupancy mask of the $i$-th dynamic agent.
+C.2 Driving Policy Evaluation
+On the NavSim benchmark, we evaluate the final agent’s performance using safety and planning metrics.
+For planning quality, we use the PDSM [dauner2024navsim], which provides a holistic score for driving quality, weighting progress, comfort, and safety.
+C.3 Baselines and Comparison Schemes
+We conduct a comprehensive comparison to validate both the world model and the downstream policy refinement.
+1. For World Model Evaluation: We compare our Impartial World Model (IWM) against strong baseline world models trained only on safe data to demonstrate the impact of optimistic bias:
+  • I2-World [liao2025i2world]: A state-of-the-art 4D occupancy world model using VQ-VAE and autoregressive transformers.
+  • DOME [gu2024dome]: A diffusion-based occupancy world model.
+2.
+For Policy Refinement: We apply our plug-and-play framework, powered by IWM, to two state-of-the-art, publicly available, pre-trained end-to-end driving agents.
+We compare the original performance against our refined versions:
+  • DiffusionDrive [liao2025diffusiondrive]: A diffusion-based planning policy.
+  • ReCogDrive [li2025recogdrive]: A VLA-based agent integrating vision-language features.
+Table 4: Reward Coefficients Configuration. Negative weights indicate penalties.
+Component Symbol Weight ($w$) Description
+Safety-Critical (Heavily Penalized)
+Collision $w_{coll}$ 20.0-20.0 Penalty if VC-IoU >0>0.
+Off-Road $w_{off}$ 10.0-10.0 Penalty for non-drivable areas.
+Clearance $w_{clr}$ 15.0-15.0 Vertical height violation.
+Comfort & Task
+Stability $w_{stab}$ 2.0-2.0 Variance of Z-axis under ego.
+Progress $w_{prog}$ +1.0+1.0 Reward for distance traveled.
+Velocity $w_{vel}$ +0.5+0.5 Alignment with speed limit.
+D Additional Experimental Results
+D.1 Ablation: Synthetic Data Ratio
+We investigate the impact of the ratio of synthetic counterfactual data mixed into the training set. Table 5 shows that a balance is crucial.
+Table 5: Effect of Synthetic Data Ratio. We find that an 80:20 mix offers the best trade-off between scene fidelity (G-IoU) and safety awareness (f-IoU).
+Real : Syn Ratio G-IoU $\uparrow$ f-IoU $\uparrow$ PDMS $\uparrow$
+100 : 0 (Baseline) 21.01 14.21 85.3
+90 : 10 35.40 38.12 88.2
+80 : 20 (Ours) 40.21 45.91 89.8
+50 : 50 38.50 46.10 87.4
+While increasing synthetic data to 50% slightly improves f-IoU (collision prediction), it degrades the general scene generation quality (G-IoU) because the synthetic physics are simplified approximations.
+The 80:20 ratio provides the optimal balance.
+D.2 Ablation: Reward Function Components
+To validate the contribution of each component in our multi-faceted reward function $R_{total}$, we conduct an ablation study based on DiffusionDrive by selectively removing specific reward terms during the RL training phase.
+Table 6 summarizes the impact on the final policy performance.
+Table 6: Ablation on Reward Components.
+Each component plays a distinct role: Safety penalties are crucial for collision avoidance, Progress rewards prevent the ”frozen robot” problem, and Comfort rewards ensure smooth 3D trajectory planning.
+Configuration Coll. Rate $\downarrow$ PDMS $\uparrow$
+DiffusionDrive 1.8% 88.1
+Full AD-R1 (Ours) 1.6% 89.8
+w/o Safety ($w_{coll},w_{off},w_{clr}=0$) 1.8% 88.5
+w/o Comfort ($w_{stab}=0$) 1.8% 87.6
+w/o Progress ($w_{prog},w_{vel}=0$) 2.0% 76.8
+Impact of Safety Penalties: When safety penalties are removed (Row 3), the Collision Rate reverts to the baseline level of 1.8%.
+While the absolute difference (0.2%) appears small, it represents a meaningful improvement given that the baseline policy is already highly optimized.
+More importantly, our Impartial World Model provides a mechanism to penalize subtle unsafe behaviors that heuristics miss.
+Impact of Comfort: Removing the 3D stability reward (Row 4) results in a PDMS of 87.6, which is even lower than the original baseline (88.1).
+This suggests that without comfort constraints, the RL process may optimize for safety or speed at the expense of ride smoothness, generating jerky trajectories that degrade the overall driving quality.
+Impact of Progress: Removing task-oriented rewards (Row 5) causes a catastrophic drop in PDMS to 76.8.
+Interestingly, the collision rate slightly increases to 2.0%, implying that an agent lacking goal-directed behavior may behave unpredictably or fail to exit dangerous zones efficiently, thereby increasing risk.
+E Qualitative Visualization
+In Figure 8, we show the behavior of a baseline agent versus our refined agent in a critical scenario.
+The original agent, failing to anticipate a sudden cut-in, generates a plan that leads to a collision.
+Our refined agent, having learned from countless imagined failures, exhibits more defensive behavior, correctly predicting the risk and executing a safe braking maneuver.
+Figure 8: Behavior of an agent with and without AD-R1 refinement. Left: The original agent’s plan results in a collision or off-road. Right: Our refined agent safely avoids the hazard.